Haste Enterprises issues 20-year, $1,000,000 bonds that pay semiannual interest of $40,000. If the effective annual rate of interest is 10%, what is the issue price of the bonds? Some relevant and irrelevant present value factors:

* PV of ordinary annuity of $1: n = 20; i = 10% is 8.51356
**PV of $1: n = 20; i = 10% is 0.14864
* PV of ordinary annuity of $1: n = 40; i = 5% is 17.5909
**PV of $1: n = 40; i = 5% is 0.14205

Multiple Choice

  • $828,000.

  • $893,000.

  • $1,000,000.

  • $1,686,000.
Answer #1

Issue price of bond is equal to the present value of all future coupon payments and the redemption value

Semi-annual rate = 10%/2 = 5%

Number of semi-annual periods = 20*2 = 40

= 40,000*PVAF(5%, 40 periods)+ 1,000,000*PVF(5%, 40 periods)

= 40,000*17.5909 + 1000,000*0.14205

= $845,686

Hence, issue price = $845,686